Calculate the lattice energy of TlBr(s) using the following thermodynamic data (all data is in kJ/mol). Note that the data given has been perturbed, so looking up the answer is probably not a good idea.

Tl(s) ΔHsublimation = 161 kJ/mol
Tl(g) IE = 569 kJ/mol
Br-Br(g) DBr-Br = 173 kJ/mol
Br(g) EA = -345 kJ/mol
TlBr(s) ΔH°f = -193 kJ/mol

_____________Kj/mol

Do you expect this value to be larger or smaller than the lattice energy of TlCl(s)? _________
Answer #1

Given Thermodynamic Data (all in kJ/mol):

  • ΔH°sublimation (Tl): 161

  • Ionization Energy (IE) of Tl(g): 569

  • Bond Dissociation Energy (Br-Br): 173

  • Electron Affinity (EA) of Br(g): -345

  • ΔH°f (TlBr(s)): -193


Steps to Calculate Lattice Energy (U):

Use the Born-Haber Cycle for TlBr(s):

ΔHf=ΔHsub+IE+12DBr-Br+EA+U

Rearrange to solve for U:

U=ΔHf(ΔHsub+IE+12DBr-Br+EA)

Plug in Values:

U=193(161+569+12×173+(345))U=193(161+569+86.5345)U=193(471.5)=664.5kJ/mol

Answer:

  • Lattice Energy of TlBr(s)−665 kJ/mol (rounded to 3 sig figs).


  • Comparison to TlCl(s): The lattice energy of TlBr(s) is smaller (less negative) than TlCl(s) because Br⁻ is larger than Cl⁻, reducing electrostatic attraction.


answered by: anonymous